give the domain and range of the following quadratic
function:
Y=3x2+ 2x -3

Answer by stanbon(75887)

give the domain and range of the following quadratic
function:
Y=3x2+ 2x -3
-------------
Domain:: All Real Numbers
-------
Range ?
Vertex occurs at x = -b/(2a) = -2/(6) = -1/3
and f(-1/3) = 3(1/9)-(2/3)- 3 = -2/3
------------------------
Since the parabola opens "up",
Range is "All Real Numbers >= -1/3
-----------
Cheers,
Stan H.
------------